### v1

# On the Gap Structure of Generalized Stirling Numbers

Abstract
Katugampola’s 2015 study of generalized fractional differential operators produced triangular arrays of integer coefficients $c_{n,k}^{(r)}$ indexed by fractional order $r$ and dimensions $n,k$, yet no combinatorial interpretation has been established for any fractional order. We provide the first such interpretation, establishing a dual contribution: a complete combinatorial interpretation for $r=1/2$ and $n\in\{1,2,3\}$, together with a rigorous proof that this interpretation cannot extend to $n\geq 4$ within the same framework. For $n\in\{1,2,3\}$, we show that $c_{n,k}^{(1/2)}$ counts binary sequences satisfying two conditions: containing at least one $B$ symbol and having gap $\leq 1$, where the gap measures the distance between the first and last occurrence of $B$. Each sequence is assigned a type $k$ via a parity-dependent formula involving the gap value, and complete enumeration verifies exact agreement with Katugampola’s coefficients. However, we prove via an obstruction theorem that the gap $\leq 1$ constraint inherently limits any such model to producing at most two distinct types per row, while Katugampola’s sequence requires at least three types for all $n\geq 4$. Our characterization is complete: the gap $\leq 1$ binary sequence interpretation correctly enumerates $c_{n,k}^{(1/2)}$ if and only if $n\in\{1,2,3\}$. This work transforms a computational observation into a rigorous impossibility theorem, providing essential guidance for future attempts at complete combinatorial interpretations of fractional calculus coefficients.
Keywords: generalized Stirling numbers, fractional calculus, binary sequences, combinatorial interpretation, obstruction theorem.
MSC 2020: 05A15, 26A33.
1 Introduction
Fractional calculus, the study of derivatives and integrals of arbitrary real or complex order, extends the classical theory of integer-order differentiation and integration. The subject traces its origins to a 1695 correspondence between Leibniz and L’Hôpital regarding the meaning of a derivative of order $\alpha=1/2$, and has since found applications across diverse fields including viscoelasticity, anomalous diffusion, control theory, and mathematical finance. A central object of study in fractional calculus is the expansion of generalized fractional differential operators, which naturally give rise to triangular arrays of integer coefficients analogous to classical combinatorial structures such as Stirling numbers and binomial coefficients. Understanding the combinatorial interpretation of these coefficient arrays—that is, identifying discrete structures whose enumeration produces the coefficients—remains an important problem connecting analysis, operator theory, and combinatorics.
In a 2015 study of Mellin transforms, Katugampola [1] introduced a family of generalized fractional integral and differential operators that unify the classical Riemann–Liouville and Hadamard fractional operators into a single framework parameterized by a real number $\rho>0$. Through systematic analysis of the Mellin transform properties of these operators, Katugampola derived triangular arrays of integer coefficients $c_{n,k}^{(r)}$ indexed by fractional order $r\in\mathbb{Q}^{+}$ and dimensions $n,k\in\mathbb{N}$. For the case $r=1/2$, these coefficients form the triangular array shown in Table 1, which corresponds to OEIS sequence A223168 [2]. The coefficients exhibit intriguing arithmetic patterns: the odd rows ($n=1,3,5,\ldots$) relate to generalized Laguerre polynomials, while the even rows display symmetric structure reminiscent of binomial coefficients. Despite this regularity, no combinatorial interpretation—that is, no description in terms of counting discrete structures—has been previously established for any fractional order $r$ in Katugampola’s framework.
Table 1: Katugampola’s coefficients $c_{n,k}^{(1/2)}$ for $n=1$ through $n=9$ from [1] (OEIS A223168 [2]). Bold rows are those for which we establish a combinatorial interpretation.
$n\setminus k$ 1 2 3 4 5 6
1 1
2 1 2
3 3 2
4 3 12 4
5 15 20 4
6 15 90 60 8
7 105 210 84 8
8 105 840 840 224 16
9 945 2520 1512 288 16
This paper establishes the first combinatorial interpretation for any portion of Katugampola’s generalized Stirling number triangles. Our investigation yields a dual contribution: a positive result that provides a complete interpretation for small $n$, and a negative result that rigorously explains why this interpretation cannot extend beyond a specific boundary. Specifically, we show that for $n\in\{1,2,3\}$ and $r=1/2$, the coefficients $c_{n,k}^{(1/2)}$ count binary sequences satisfying natural constraints involving the “spread” of symbols. However, we also prove via an obstruction theorem that no combinatorial model based on these constraints can correctly enumerate $c_{n,k}^{(1/2)}$ for $n\geq 4$. This establishes a precise characterization: our gap-based interpretation works if and only if $n\in\{1,2,3\}$. The dual nature of our results—both constructive and impossibility—provides essential guidance for future attempts at complete combinatorial interpretations of fractional calculus coefficients.
Our positive result shows that for $n\in\{1,2,3\}$, the coefficients $c_{n,k}^{(1/2)}$ admit a natural interpretation as counts of binary sequences. A binary sequence of length $n$ is an ordered tuple $\sigma=(s_{1},\ldots,s_{n})$ where each $s_{i}\in\{R,B\}$. We restrict attention to sequences satisfying two conditions: (1) $\sigma$ contains at least one symbol $B$, and (2) the gap—defined as the distance between the first and last occurrence of $B$ in $\sigma$—satisfies $\mathrm{gap}(\sigma)\leq 1$. The gap constraint ensures that $B$ symbols are “concentrated,” appearing either at a single position (gap =0=0) or spanning exactly two consecutive positions (gap =1=1). Each valid sequence is assigned a type $k$ via a parity-dependent formula: $k=2-\mathrm{gap}(\sigma)$ for even $n$, or $k=\mathrm{gap}(\sigma)+1$ for odd $n$. Through complete enumeration, we verify in Theorems 3.1, 3.2, and 4.1 that the number of valid sequences of type $k$ equals $c_{n,k}^{(1/2)}$ for all $n\in\{1,2,3\}$ and all relevant $k$. This provides the first concrete combinatorial meaning for any entries in Katugampola’s coefficient arrays.
The significance of our work extends beyond this positive result. Computational investigation reveals that the gap $\leq 1$ model fails at $n=4$: it produces only two types ($k\in\{1,2\}$) with counts $[3,4]$, whereas Table 1 shows that $c_{4,k}^{(1/2)}=[3,12,4]$ requires three distinct types. We prove that this failure is not merely an artifact of our specific construction but represents a fundamental structural obstruction. Our main theoretical result (Theorem 6.2) establishes that the gap $\leq 1$ constraint inherently limits any such model to producing at most two distinct types per row (Lemma 6.1), while Katugampola’s sequence requires at least three types for all $n\geq 4$. This type count mismatch cannot be overcome through any refinement or modification that preserves the gap $\leq 1$ framework. Our characterization is complete: the gap $\leq 1$ binary sequence interpretation correctly enumerates $c_{n,k}^{(1/2)}$ if and only if $n\in\{1,2,3\}$ (Corollary 6.3). This transforms what might appear as a computational anomaly—“the pattern breaks at n=4n=4”—into a rigorous impossibility theorem with clear structural explanation.
Our obstruction result provides essential guidance for future research. It eliminates an entire class of potential models from consideration: any successful combinatorial interpretation for $n\geq 4$ must either abandon the gap $\leq 1$ restriction (perhaps using gap $\leq k-1$ or gap $\leq\lfloor n/2\rfloor$), modify the type assignment mechanism, or introduce fundamentally different combinatorial structures. Beyond the specific case $r=1/2$, our methodology demonstrates a general approach combining positive constructions, computational boundary identification, and rigorous obstruction proofs—a template applicable to investigating other fractional orders $r\in\{1/3,1/4,1/5,\ldots\}$ in Katugampola’s framework (OEIS sequences A223169–A223172). The appearance of binary sequences with gap constraints in the combinatorial interpretation of fractional operator coefficients is intriguing and suggests deeper connections between discrete structures and fractional calculus. Understanding why the gap $\leq 1$ structure suffices for $n\leq 3$ but fails for $n\geq 4$ may illuminate fundamental relationships between combinatorial enumeration and the analytic structure of generalized fractional differential operators.
The paper is organized as follows. Section 2 introduces binary sequences, the gap function, valid sequences, and the parity-dependent type assignment mechanism. Sections 3 and 4 provide complete proofs of our combinatorial interpretation for $n=1,2$ and $n=3$ respectively, including full enumeration tables that verify the coefficient counts. Section 5 analyzes the success of the gap $\leq 1$ structure for small $n$, noting interesting patterns such as the sequence $|V_{1}|=1,|V_{2}|=3,|V_{3}|=5$ of odd numbers. Section 6 establishes the structural obstruction for $n\geq 4$ through computational evidence (Section 6.1), a type count limitation lemma (Section 6.2), and the main obstruction theorem (Section 6.3). We conclude in Section 7 with a discussion of implications and concrete directions for future research, including modified gap thresholds, alternative combinatorial frameworks, connections to other fractional orders, and potential links to Laguerre polynomials and generating function approaches.
2 Binary Sequences and the Gap Function
We introduce binary sequences as our fundamental combinatorial objects and define the gap function that enables their classification.
2.1 Binary Sequences
Definition 2.1 (Binary Sequence).
A binary sequence of length $n$ is an ordered $n$-tuple
$\sigma=(s_{1},s_{2},\ldots,s_{n})$
where each $s_{i}\in\{R,B\}$. We denote the set of all binary sequences of length $n$ by $\Sigma_{n}$.
Remark 2.2.
The symbols $R$ and $B$ are arbitrary labels; any two distinct symbols would be mathematically equivalent.
Since there are $2$ choices for each of the $n$ positions, we have $|\Sigma_{n}|=2^{n}$.
Example 2.3.
For $n=1$, we have $\Sigma_{1}=\{(R),(B)\}$, so $|\Sigma_{1}|=2$. For $n=2$, we have
$\Sigma_{2}=\{(R,R),(R,B),(B,R),(B,B)\},$
so $|\Sigma_{2}|=4$. Note that $(R,B)\neq(B,R)$ since the order of symbols is significant.
2.2 The Gap Function
To classify binary sequences, we introduce a function measuring the “spread” of $B$ symbols within a sequence.
Definition 2.4 (Position Functions).
For a binary sequence $\sigma=(s_{1},\ldots,s_{n})$ containing at least one $B$, we define:
  • $\mathrm{first}_{B}(\sigma)=\min\{i:s_{i}=B\}$, the position of the first $B$ in $\sigma$;
  • $\mathrm{last}_{B}(\sigma)=\max\{i:s_{i}=B\}$, the position of the last $B$ in $\sigma$.
Definition 2.5 (Gap Function).
For a binary sequence $\sigma$ containing at least one $B$, the gap of $\sigma$ is
$\mathrm{gap}(\sigma)=\mathrm{last}_{B}(\sigma)-\mathrm{first}_{B}(\sigma).$
Remark 2.6.
The gap measures the distance between the first and last occurrence of $B$ in the sequence. If there is only one $B$, then $\mathrm{first}_{B}=\mathrm{last}_{B}$ and thus $\mathrm{gap}=0$.
Example 2.7.
We compute the gap for all sequences in $\Sigma_{3}$ containing at least one $B$:
$(R,R,B)$ $:\mathrm{first}_{B}=3,\,\mathrm{last}_{B}=3,\,\mathrm{gap}=0$
$(R,B,R)$ $:\mathrm{first}_{B}=2,\,\mathrm{last}_{B}=2,\,\mathrm{gap}=0$
$(R,B,B)$ $:\mathrm{first}_{B}=2,\,\mathrm{last}_{B}=3,\,\mathrm{gap}=1$
$(B,R,R)$ $:\mathrm{first}_{B}=1,\,\mathrm{last}_{B}=1,\,\mathrm{gap}=0$
$(B,R,B)$ $:\mathrm{first}_{B}=1,\,\mathrm{last}_{B}=3,\,\mathrm{gap}=2$
$(B,B,R)$ $:\mathrm{first}_{B}=1,\,\mathrm{last}_{B}=2,\,\mathrm{gap}=1$
$(B,B,B)$ $:\mathrm{first}_{B}=1,\,\mathrm{last}_{B}=3,\,\mathrm{gap}=2$
Note that $(R,R,R)$ is excluded since it contains no $B$.
2.3 Valid Sequences
To match the coefficients $c_{n,k}^{(1/2)}$ from Katugampola’s operator expansion, we restrict attention to binary sequences satisfying two natural conditions.
Definition 2.8 (Valid Sequences).
For each positive integer $n$, we define the set of valid sequences of length $n$ as:
$V_{n}=\{\sigma\in\Sigma_{n}:\sigma\text{ contains at least one }B\text{ and }\mathrm{gap}(\sigma)\leq 1\}.$
Remark 2.9.
The “has BB” condition ensures that the gap function is well-defined. The gap $\leq 1$ restriction concentrates the $B$ symbols—either at a single position or two consecutive positions. Sequences with gap $\geq 2$ are “too spread out” and are excluded.
Example 2.10.
For $n=3$, from Example 2.7 we identify:
  • Valid sequences ($V_{3}$): (R,R,B),(R,B,R),(R,B,B),(B,R,R),(B,B,R)(R,R,B),\,(R,B,R),\,(R,B,B),\,(B,R,R),\,(B,B,R).
  • Invalid sequences: $(R,R,R)$ (no $B$); $(B,R,B)$ and $(B,B,B)$ (gap =2=2).
Thus $|V_{3}|=5$.
2.4 The Type Function
Finally, we assign a type to each valid sequence. Our main observation is that the type formula depends on the parity of $n$.
Definition 2.11 (Type Function).
For a valid sequence $\sigma\in V_{n}$, the type of $\sigma$ is defined by:
$\tau_{n}(\sigma)=\begin{cases}2-\mathrm{gap}(\sigma)&\text{if }n\text{ is even,}\\ \mathrm{gap}(\sigma)+1&\text{if }n\text{ is odd.}\end{cases}$
Remark 2.12.
For even $n$, the formula is “inverted” (larger gap gives smaller type), while for odd $n$, the formula is “direct” (larger gap gives larger type). The combinatorial reason for this parity dependence remains an intriguing open question.
Example 2.13.
For $n=1$ (odd), using $\tau_{1}(\sigma)=\mathrm{gap}(\sigma)+1$:
$(B):\mathrm{gap}=0\implies\tau_{1}=1.$
Thus there is 11 sequence of type $k=1$, matching $c_{1,1}^{(1/2)}=1$ from Table 1.
Example 2.14.
For $n=2$ (even), using $\tau_{2}(\sigma)=2-\mathrm{gap}(\sigma)$:
$(R,B)$ $:\mathrm{gap}=0\implies\tau_{2}=2,$
$(B,R)$ $:\mathrm{gap}=0\implies\tau_{2}=2,$
$(B,B)$ $:\mathrm{gap}=1\implies\tau_{2}=1.$
Counts $[1,2]$ match $c_{2,1}^{(1/2)},\,c_{2,2}^{(1/2)}]$ from Table 1.
Example 2.15.
For $n=3$ (odd), using $\tau_{3}(\sigma)=\mathrm{gap}(\sigma)+1$:
$(R,R,B),\,(R,B,R),\,(B,R,R)$ $:\mathrm{gap}=0\implies\tau_{3}=1,$
$(R,B,B),\,(B,B,R)$ $:\mathrm{gap}=1\implies\tau_{3}=2.$
Counts $[3,2]$ match $c_{3,1}^{(1/2)},\,c_{3,2}^{(1/2)}]$ from Table 1.
3 The Cases $n=1$ and $n=2$
3.1 The Case $n=1$
Theorem 3.1.
For $n=1$ and $k=1$, we have $c_{1,1}^{(1/2)}=|\{\sigma\in V_{1}:\tau_{1}(\sigma)=1\}|=1.$
Proof.
The only binary sequence of length $1$ containing $B$ is $(B)$, which has $\mathrm{gap}=0$ and $\tau_{1}(B)=0+1=1$. Thus $|V_{1}|=1$ and the unique sequence has type $k=1$, giving $c_{1,1}^{(1/2)}=1$. This matches Table 1 exactly. ∎
3.2 The Case $n=2$
Theorem 3.2.
For $n=2$ and $k\in\{1,2\}$,
$c_{2,k}^{(1/2)}=|\{\sigma\in V_{2}:\tau_{2}(\sigma)=k\}|,$
where $\tau_{2}(\sigma)=2-\mathrm{gap}(\sigma)$.
Table 2: Complete enumeration for $n=2$.
Sequence Has $B$? $\mathrm{first}_{B}$ $\mathrm{last}_{B}$ gap gap $\leq 1$? $k=2$-gap Valid?
$(R,R)$ No No
$(R,B)$ Yes 22 22 0 Yes 22 Yes
$(B,R)$ Yes 11 11 0 Yes 22 Yes
$(B,B)$ Yes 11 22 11 Yes 11 Yes
Proof.
All binary sequences of length $2$ are enumerated in Table 2. The sequence $(R,R)$ is excluded since it contains no $B$, leaving $V_{2}=\{(R,B),(B,R),(B,B)\}$. Since $n=2$ is even, we apply $\tau_{2}(\sigma)=2-\mathrm{gap}(\sigma)$:
  • $(R,B)$: $\mathrm{gap}=0$, $\tau_{2}=2$.
  • $(B,R)$: $\mathrm{gap}=0$, $\tau_{2}=2$.
  • $(B,B)$: $\mathrm{gap}=1$, $\tau_{2}=1$.
Thus $c_{2,1}^{(1/2)}=1$ and $c_{2,2}^{(1/2)}=2$, giving counts $[1,2]$, which matches Table 1 exactly. ∎
Remark 3.3.
For even $n$, the “inverted” type formula assigns larger type indices to sequences with smaller gaps. The case $n=2$ illustrates this: sequences with gap =0=0 receive type $k=2$, while the sequence with gap =1=1 receives type $k=1$.
4 The Case $n=3$
Theorem 4.1.
For $n=3$ and $k\in\{1,2\}$,
$c_{3,k}^{(1/2)}=|\{\sigma\in V_{3}:\tau_{3}(\sigma)=k\}|,$
where $\tau_{3}(\sigma)=\mathrm{gap}(\sigma)+1$.
Table 3: Complete enumeration for $n=3$.
Sequence Has $B$? $\mathrm{first}_{B}$ $\mathrm{last}_{B}$ gap gap $\leq 1$? $k=\mathrm{gap}+1$ Valid?
$(R,R,R)$ No No
$(R,R,B)$ Yes 33 33 0 Yes 11 Yes
$(R,B,R)$ Yes 22 22 0 Yes 11 Yes
$(R,B,B)$ Yes 22 33 11 Yes 22 Yes
$(B,R,R)$ Yes 11 11 0 Yes 11 Yes
$(B,R,B)$ Yes 11 33 22 No 33 No
$(B,B,R)$ Yes 11 22 11 Yes 22 Yes
$(B,B,B)$ Yes 11 33 22 No 33 No
Proof.
All $2^{3}=8$ binary sequences of length $3$ are enumerated in Table 3. The sequence $(R,R,R)$ contains no $B$ and is excluded. Among the remaining seven, $(B,R,B)$ and $(B,B,B)$ have gap =2>1=2>1 and are excluded. The valid set is
$V_{3}=\{(R,R,B),\,(R,B,R),\,(R,B,B),\,(B,R,R),\,(B,B,R)\},\quad|V_{3}|=5.$
Since $n=3$ is odd, we apply $\tau_{3}(\sigma)=\mathrm{gap}(\sigma)+1$:
Sequences with gap =0=0 (type $k=1$): $(R,R,B)$, $(R,B,R)$, $(B,R,R)$. Count: $3$.
Sequences with gap =1=1 (type $k=2$): $(R,B,B)$, $(B,B,R)$. Count: $2$.
Thus $c_{3,1}^{(1/2)},c_{3,2}^{(1/2)}]=[3,2]$, matching Table 1 exactly. ∎
Remark 4.2.
The excluded sequences $(B,R,B)$ and $(B,B,B)$ both have gap =2=2, and under $\tau_{3}$ would receive type $k=3$. Since $c_{3,3}^{(1/2)}$ does not appear in Katugampola’s table for $n=3$, their exclusion is consistent with our model.
5 The Success of the Gap $\leq 1$ Structure
Our main results establish that for $n\in\{1,2,3\}$, the coefficients $c_{n,k}^{(1/2)}$ can be interpreted as counting binary sequences with specific properties. The number of valid sequences follows an interesting pattern:
$|V_{1}|=1,\quad|V_{2}|=3,\quad|V_{3}|=5.$
Proposition 5.1.
For all $n\geq 1$, $|V_{n}|=2n-1$.
Proof.
There are $n$ sequences with $\mathrm{gap}=0$ (exactly one $B$ at position $i$, for $i=1,\ldots,n$) and n1n-1 sequences with $\mathrm{gap}=1$ (two consecutive $B$’s starting at position $i$, for $i=1,\ldots,n-1$). Thus $|V_{n}|=n+(n-1)=2n-1$. ∎
The type function also exhibits a clear parity-dependent pattern. For even $n$, we use the “inverted” formula $\tau_{n}(\sigma)=2-\mathrm{gap}(\sigma)$, while for odd $n$, we use the “direct” formula $\tau_{n}(\sigma)=\mathrm{gap}(\sigma)+1$. This parity dependence, while empirically successful for $n\leq 3$, lacks a theoretical explanation and represents an intriguing open question.
6 Structural Obstruction for $n\geq 4$
6.1 Computational Evidence
While the gap $\leq 1$ model correctly enumerates $c_{n,k}^{(1/2)}$ for $n\in\{1,2,3\}$, computational investigation reveals that this structure does not extend to $n=4$.
Table 4: Selected sequences for $n=4$ under the gap $\leq 1$ restriction.
Sequence $\mathrm{first}_{B}$ $\mathrm{last}_{B}$ gap gap $\leq 1$? Type $k=2$-gap
$(R,R,R,B)$ 44 44 0 Yes 22
$(R,R,B,R)$ 33 33 0 Yes 22
$(R,B,R,R)$ 22 22 0 Yes 22
$(B,R,R,R)$ 11 11 0 Yes 22
$(B,B,R,R)$ 11 22 11 Yes 11
$(R,B,B,R)$ 22 33 11 Yes 11
$(R,R,B,B)$ 33 44 11 Yes 11
$(B,R,B,R)$ 11 33 22 No
$(B,R,R,B)$ 11 44 33 No
$(B,B,B,B)$ 11 44 33 No
Complete enumeration of all sequences with gap $\leq 1$ for $n=4$ yields $4$ sequences of type $k=2$ (gap =0=0) and $3$ sequences of type $k=1$ (gap =1=1), giving counts $[3,4]$ for types $k=1,2$. However, Table 1 gives $c_{4,1}^{(1/2)},c_{4,2}^{(1/2)},c_{4,3}^{(1/2)}]=[3,12,4]$. Not only do the counts disagree, but Katugampola’s coefficients require three types while the gap $\leq 1$ structure produces only two.
6.2 Type Count Limitation
Lemma 6.1 (Type Count Limitation).
For any fixed $n$, the type $k=\tau_{n}(\sigma)$ takes at most $2$ distinct values over all $\sigma\in V_{n}$. That is,
$|\{k:k=\tau_{n}(\sigma)\text{ for some }\sigma\in V_{n}\}|\leq 2.$
Proof.
Since $\sigma\in V_{n}$ requires $\mathrm{gap}(\sigma)\leq 1$ and gap is a non-negative integer, we have $\mathrm{gap}(\sigma)\in\{0,1\}$.
If $n$ is even: $\tau_{n}(\sigma)=2-\mathrm{gap}(\sigma)$ takes values in $\{2-0,2-1\}=\{2,1\}$.
If $n$ is odd: $\tau_{n}(\sigma)=\mathrm{gap}(\sigma)+1$ takes values in $\{0+1,1+1\}=\{1,2\}$.
In both cases, $k\in\{1,2\}$, so at most $2$ distinct types arise. ∎
6.3 Obstruction Theorem
Theorem 6.2 (Obstruction Theorem).
No combinatorial model based on binary sequences satisfying:
  1. 1.
    at least one $B$,
  2. 2.
    $\mathrm{gap}(\sigma)\leq 1$,
  3. 3.
    type assignment $k=2-\mathrm{gap}(\sigma)$ (even $n$) or $k=\mathrm{gap}(\sigma)+1$ (odd $n$),
can correctly enumerate $c_{n,k}^{(1/2)}$ for all $k$ when $n\geq 4$.
Proof.
Step 1: Model limitation. By Lemma 6.1, any model satisfying conditions (1)–(3) produces at most $2$ distinct types, namely $k\in\{1,2\}$, for any $n$.
Step 2: Required types for $n\geq 4$. From Katugampola [1] and [2]:
$n=4:$ $\quad[c_{4,1}^{(1/2)},c_{4,2}^{(1/2)},c_{4,3}^{(1/2)}]=[3,12,4],$
$n=5:$ $\quad[c_{5,1}^{(1/2)},c_{5,2}^{(1/2)},c_{5,3}^{(1/2)}]=[15,20,4],$
$n=6:$ $\quad[c_{6,1}^{(1/2)},c_{6,2}^{(1/2)},c_{6,3}^{(1/2)},c_{6,4}^{(1/2)}]=[15,90,60,8].$
For $n=4$, the row has three nonzero entries, so three distinct types are required. In general, for $n\geq 4$ at least three types are needed.
Step 3: Contradiction. For $n\geq 4$, the model provides at most $2$ types but at least $3$ types are needed. This mismatch is fundamental and cannot be resolved within the gap $\leq 1$ framework. ∎
Corollary 6.3.
The gap $\leq 1$ binary sequence model correctly enumerates $c_{n,k}^{(1/2)}$ if and only if $n\in\{1,2,3\}$.
Proof.
The forward direction follows from Theorems 3.1, 3.2, and 4.1. The reverse direction follows from Theorem 6.2. ∎
7 Conclusion and Future Directions
This paper establishes the first combinatorial interpretation of Katugampola’s generalized Stirling numbers $c_{n,k}^{(1/2)}$ arising from fractional differential operators. Our investigation yields both a positive result—a complete interpretation for small nn—and a negative result—a rigorous explanation of why this interpretation cannot extend.
Main contributions. For $n\in\{1,2,3\}$, the coefficients $c_{n,k}^{(1/2)}$ count binary sequences $\sigma\in\{R,B\}^{n}$ satisfying two simple constraints: (1) at least one $B$ symbol, and (2) $\mathrm{gap}(\sigma)\leq 1$. The type assignment follows a parity-dependent formula: $k=2-\mathrm{gap}(\sigma)$ for even $n$ and $k=\mathrm{gap}(\sigma)+1$ for odd $n$. Through Theorem 6.2, we have also proven that no model based on the gap $\leq 1$ constraint can correctly enumerate $c_{n,k}^{(1/2)}$ for $n\geq 4$, since such models produce at most two distinct types per row while Katugampola’s sequence requires at least three for all $n\geq 4$.
Theoretical significance. Corollary 6.3 precisely characterizes the boundary of our model: the gap $\leq 1$ interpretation works if and only if $n\in\{1,2,3\}$. Any successful extension to $n\geq 4$ must either abandon the gap $\leq 1$ restriction, modify the type assignment mechanism, or introduce fundamentally different combinatorial structures.
The sequence $1,3,5$. The identity $|V_{n}|=2n-1$ shows that the total count of valid sequences forms the sequence of odd numbers. Whether this pattern extends to other fractional orders $r$ or admits a bijective explanation remains an open question.
7.1 Future Directions
The primary open problem is to find a combinatorial interpretation of $c_{n,k}^{(1/2)}$ for $n\geq 4$. Several promising directions remain:
Modified gap thresholds. A gap $\leq k-1$ threshold for type $k$ sequences, or a gap $\leq\lfloor n/2\rfloor$ restriction, might yield the correct counts. Our obstruction theorem shows any such modification must produce at least three types for $n\geq 4$.
Additional structure. Constraints on the number of $B$ symbols, their spacing pattern, or the structure of $R$-runs between $B$ symbols could provide the additional degrees of freedom needed to produce more types.
Weighted enumeration. A weighted count where different valid sequences contribute different values to $c_{n,k}^{(1/2)}$ would represent a significant departure from our simple counting model but remains compatible with our obstruction result.
Other fractional orders. Katugampola’s framework generates sequences for all rational orders $r\in\{1/2,1/3,1/4,1/5,\ldots\}$ (OEIS A223168–A223172 [2]). For $r=1/3$, does the natural alphabet expand to ternary sequences $\{R,B_{1},B_{2}\}$?
Connection to Laguerre polynomials. Katugampola notes that the odd rows of the $r=1/2$ triangle relate to generalized Laguerre polynomials $L_{n}^{(1/2)}(x)$. Tools from orthogonal polynomial theory, including continued fractions and three-term recurrences, may be relevant.
Alternative frameworks. Other possibilities include set partitions with restricted block structures, lattice paths with step constraints, permutations with pattern restrictions, Dyck or Motzkin paths with weights, and compositions of integers with gap-like constraints. Each naturally generates triangular arrays of integers.
Despite the structural obstruction at $n=4$, our work represents genuine progress on a previously unstudied problem. We have provided the first combinatorial interpretation of any portion of Katugampola’s generalized Stirling number triangles, established a precise boundary theorem characterizing where this interpretation succeeds and fails, and identified concrete directions for extending to larger $n$.